It has been observed that some persons who suffer colitis, again suffer colitis within one year of the first episode. This is due, in part, to damage from the first episode. The performance of a new drug designed to prevent a second episode is to be tested for its effectiveness in preventing a second episode. In order to do this two groups of people suffering a first episode are selected. There are 75 people in the first group and this group will be administered the new drug. There are 45 people in the second group and this group will be administered a placebo. After one year, 11% of the first group has a second episode and 9% of the second group has a second episode. Conduct a hypothesis test to determine, at the significance level 0.05, whether there is reason to believe that the true percentage of those in the first group who suffer a second episode is less than the true percentage of those in the second group who suffer a second episode? Select the [Rejection Region, Decision to Reject (RH0) or Failure to Reject (FRH0)].
Answer #1

H0: The true percentage of those in the first group who suffer a second episode is equal to the true percentage of those in the second group who suffer a second episode. That is p1 = p2

H1: The true percentage of those in the first group who suffer a second episode is less than the true percentage of those in the second group who suffer a second episode. That is p1 < p2

Given, observed proportions are,

= 0.11

= 0.09

Pooled proportion, p = (n1 + n2 ) / (n1 + n2) = (75 * 0.11 + 45 * 0.09) / (75 + 45) = 0.1025

Standard error of difference in proportions = SE =

= 0.0572

Test statistic, z = ( - ) / SE = (0.11 - 0.09)/0.0572 = 0.3497

P-value = P(z < 0.3497) = 0.6367

Since, p-value is greater than 0.05 significance level, we fail to reject null hypothesis H0 and conclude that there is no strong evidence that the true percentage of those in the first group who suffer a second episode is less than the true percentage of those in the second group who suffer a second episode.

Failure to Reject (FRH0)]